# Why the Agent Made that Decision: Explaining Deep Reinforcement Learning with Vision Masks

### v1

@@ Abstract
Due to the inherent lack of transparency in deep neural networks, it is challenging for deep reinforcement learning (DRL) agents to gain trust and acceptance from users, especially in safety-critical applications such as medical diagnosis and military operations. Existing methods for explaining an agent’s decision either require to retrain the agent using models that support explanation generation or rely on perturbation-based techniques to reveal the significance of different input features in the decision making process. However, retraining the agent may compromise its integrity and performance, while perturbation-based methods have limited performance and lack knowledge accumulation or learning capabilities. Moreover, since each perturbation is performed independently, the joint state of the perturbed inputs may not be physically meaningful. To address these challenges, we introduce VisionMask, a standalone explanation model trained end-to-end to identify the most critical regions in the agent’s visual input that can explain its actions. VisionMask is trained in a self-supervised manner without relying on human-generated labels. Importantly, its training does not alter the agent model, hence preserving the agent’s performance and integrity. We evaluate VisionMask on Super Mario Bros (SMB) and three Atari games. Compared to existing methods, VisionMask achieves a 14.9% higher insertion accuracy and a 30.08% higher F1-Score in reproducing original actions from the selected visual explanations. We also present examples illustrating how VisionMask can be used for counterfactual analysis.
@@ 1 Introduction
Original State LIME RISE Greydanus et al. SARFA VisionMask
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Visual comparison of saliency maps generated by VisionMask, LIME(Ribeiro, Singh, and Guestrin 2016), RISE(Petsiuk, Das, and Saenko 2018), Greydanus et al. (2018), and SARFA(Puri et al. 2020) for the PPO agent(Schulman et al. 2017; Nguyen 2020) trained in the Super Mario Bros.(Kauten 2018) environment. Saliency level ranges from blue(low) to red(high). Row 1: The chosen action is to move left. The explanation is that Mario is attempting to avoid the Piranha Plants on top of the green pipes. VisionMask accurately identified the two Piranha Plants on the pipes. While RISE and Greydanus highlighted almost the entire image and LIME and SARFA highlighted areas not directly related to the explanation. Row 2:The chosen action is to jump right with acceleration. The explanation is that Mario intends to land on top of the Koopas to clear his path and get reward. Again VisionMask clearly highlighted the two Koopas whereas the other saliency maps failed to identify the correct region.
Deep Reinforcement Learning (DRL) is a powerful technology in machine intelligence, widely used for many applications, including ChatGPT (Brown et al. 2020; Achiam et al. 2023) and Alpha Go Zero (Silver et al. 2017). It enables effective end-to-end training of decision models that map an agent’s observations to actions, aiming to maximizing the long term expected rewards. However, understanding a DRL agent’s decision-making process is challenging, due to the inherent lack of explainability in the high-dimensional, non-linear structure of its underlying Deep Neural Network (DNN). The lack of transparency undermines human users’ trust and acceptance, particularly in safety-critical scenarios, such as medical diagnosis and military operations. To ensure wide acceptance and efficient use of this technology, it is crucial to provide accurate and human-interpretable explanations. These considerations motivated the recent interests in Explainable Reinforcement Learning (XRL).
Some existing works integrate explanation generation into the agent model and train them together. We refer to these as model-based approaches. For example, Interpretable DQN (i-DQN) (Annasamy and Sycara 2019) implements a deep Q-network with an attention mechanism, using the attention weights between state and action embeddings to represent their importance. Saliency-guided Q-networks (SGQNs) (Bertoin et al. 2022) calculate the importance of inputs as a derivative of the Q-function and train the critic model to both minimize the value prediction error and maximize the fidelity of the explanation. Some other works moved away from DNN and adopt special policy architectures that are inherently more interpretable (Liu et al. 2018; Verma et al. 2018; Bastani, Pu, and Solar-Lezama 2018). However, these models usually experience performance loss compared to the DNN-based policy model due to the limitations imposed by the special structure. Those model-based approaches cannot provide explanations for a trained DRL agent, unless they retrain the agent using their specific models, which is prohibitively expensive. And the retrain process may compromise the agent’s performance and integrity as it strives to maintain a balance between the optimality and explainability.
One of the most widely used techniques to provide explanations for machine learning models is the attribution-based method, which assigns credits to input features based on their contribution to AI agent’s decisions. This technique was originally developed to explain image classification models in computer vision. It generates saliency maps by perturbing the input image to reveal their contributions to the classification process without modifying the classifier itself (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016). Inspired by this, several works have explored to use attribution-based technique to explain DRL models while treating the agent as a black-box. Greydanus et al. (2018) perturb the input features using Gaussian blur at fixed intervals and measure how the perturbation changes the action probability. The difference in action probability before and after perturbation are used to generate a saliency-map for the importance of each input feature. Similarly, Iyer et al. (2018) and Puri et al. (2020) compute the importance score based on the differences in Q values before and after the perturbation. The latter also improve the accuracy of the saliency map by normalizing the effects of perturbation between target and non-target actions.
The perturbation-based methods do not alter the agent model, nor do they impose any specific requirements on it. Hence they are applicable to any trained DRL agents. However, these techniques have some significant limitations. First, to manage complexity, perturbations are performed according to a set of predefined rules, such as using Gaussian blur or occluding with a constant value, which severely limits the search space they can explore. For DRL agents that map visual input directly to actions, due to the large input dimensions, there is no guarantee that the predefined perturbation can find the set of pixels that trigger the agent decision. Secondly, perturbation-based methods do not learn from past experiences, nor do they improve performance by comparing similarities and differences across different input states and actions. Thirdly, each input feature (e.g., input pixel) is perturbed independently, with no guarantee that the joint state of the perturbed pixels represents meaningful objects. As shown in Table 1, the saliency maps generated by the perturbation-based method either consider all pixels equally important or identify regions in the image that do not contain meaning objects.
To address the aforementioned challenges, we propose VisionMask, a model trained end-to-end to generate attribution-based explanations for vision-based DRL without perturbation or altering the agent. VisionMask ouputs a saliency map that highlights the critical regions in the agent’s visual input responsible for triggering the agent’s action. The saliency map reveals the importance of each pixel in the decision making process, and will also be referred to as importance map or importance mask in the rest of the paper. VisionMask is trained using self-supervised contrastive learning. No human generated label is required. The training of VisionMask does not alter the agent model, thereby preserving the agent’s performance and integrity. The (fixed) policy model of the DRL agent is integrated into the training pipeline to generate the loss, and involved in the backpropagation to pass the gradient to the mask generation model. Therefore, each VisionMask model is trained to explain a specific DRL agent. The training process is agent agnostic, meaning that it works with any agent with a policy model that maps visual input to actions. VisionMasks’s learnable capability allows it to accumulate knowledge from previous experiences to improve future explanations. During testing time, VisionMask does not need to access agent’s internal state to generate the importance mask. The mask is generated by a forward pass of the VisionMask model.
As shown in Table 1, VisionMask accurately identifies the critical regions in the visual input that explains the agent’s action. Beyond providing precise explanations directly from the visual input, VisionMask also facilitates counterfactual analysis. By selectively activating or deactivating different critical regions, meaningful counterfactual scenarios can be generated and assessed. This is similar to the perturbation-based approach, however, at a much higher level.
@@
  • We present VisionMask, a novel attribution method for explainable visual reinforcement learning that generates action-specific importance masks. Generated using a single forward pass, the mask does not only highlight the regions in the visual input that explain the agent’s behavior but also facilitates counterfactual analysis.
  • VisionMask is model-agnostic in that it works with any vision-based DRL agent without modifying the agent model.
  • VisionMask is trained using self-supervised contrastive learning. No human intervention is needed for data labeling.
  • The fidelity of the explanation is measured by the agreement between the agent’s actions generated from the original visual input versus from the masked visual input. We quantify this agreement using accuracy, precision, recall and F1-score. In average, VisionMask outperforms other attribution-based explanation models by 30.08% on F1-score and 14.9% on insertion accuracy..
@@ 2 Background and related work
The nested structure and non-linear operation of DNN make it challenging for humans to understand how the outputs are derived from inputs. Some existing works address this challenge by explaining model outputs through input attributions (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016; Fong and Vedaldi 2017). For example, Randomized Input Sampling for Explanation of Black-box Models (RISE) (Petsiuk, Das, and Saenko 2018) is a perturbation-based approach that explains neural image classifiers by applying randomly generated masks to the input image and assessing their impact. Local Interpretable Model-agnostic Explanations (LIME)(Ribeiro, Singh, and Guestrin 2016) tries to explain local instances by approximating them within a nearby vicinity using a linear decision model, where the explainable elements are superpixels (i.e., small image regions of similar pixels).
Built on top of the DNNs, the lack of transparency of the DRL agents (Samek, Wiegand, and Müller 2017; Ras et al. 2022; Angelov and Soares 2020) undermines humans trust and hinders their adoption. Several studies have addressed this challenging problems. According to (Qing et al. 2022), existing efforts can be categorized into four main approaches: model explaining, reward explaining, task explaining, and state explaining.
Model explaining relies on inherently interpretable model architectures or auxiliary reasoning mechanisms to generate explanations (Topin et al. 2021; Jiang and Luo 2019; Landajuela et al. 2021). However, these self-explanatory or symbolic models often suffer from decreased performance compared to state-of-the-art neural network based RL policies and may lack the representational power needed to learn more complex policies. Reward explanation typically involves using explicitly designed explainable reward functions to generate explanations (Juozapaitis et al. 2019; Ashwood, Jha, and Pillow 2022). Task explanation considers a policy as a composition of sub-task and action sequences, explaining the behavior in terms of relationships among sub-tasks (Shu, Xiong, and Socher 2017; Sodhani, Zhang, and Pineau 2021). These approaches often assume the existence of a reward decomposition scheme or a predefined sub-task partition mechanism, which may not hold true in all RL environments.
@@
State explaining methods can further be divided into three categories: attention-based, Shapley value-based and perturbation-based mechanisms, which are detailed below. Attention is a common approach used for explainable RL in several existing works (Peng, Riedl, and Ammanabrolu 2022; Xu et al. 2020; Annasamy and Sycara 2019). However, similar to i-DQN (Annasamy and Sycara 2019), these methods cannot explain a given DRL model because their attention model must be trained alongside the agent model. Shapley value (Shapley et al. 1953), a concept from game theory, has also been introduced into XRL. SVERL (Beechey, Smith, and Şimşek 2023) provides a theoretical analysis on how Shapley values can be applied to explain value and policy network in DRL. However, this approach is still in its early stages and assumes prior knowledge of the transition model of the environment, which is not realistic for realistic applications.
Perturbation-based methods (Greydanus et al. 2018; Iyer et al. 2018; Puri et al. 2020) compute saliency maps of the input features by comparing the action probability or value function before and after perturbation. Greydanus et al. (2018) perturbed the state with Gaussian blur at fixed intervals, while Iyer et al. (2018) identified objects and perturbed them with a constant gray value. Puri et al. (2020) improved on Greydanus et al. (2018) by removing the effect of perturbations from other irrelevant actions. As discussed in section 1, the performance of these techniques is constrained by predefined perturbation rules and the lack of knowledge accumulation. Furthermore, there is no guarantee that the perturbed input remains physically meaningful.
@@ 3 VisionMask
This section presents VisionMask, an innovative agent-agnostic framework for XRL. The primary goal is to produce action-wise importance-mask that attribute explainable elements in the current state to each possible action. The definition of explainable elements may vary across different applications. However, in this paper, we focus on identifying explanatory pixel regions in the agent’s visual input.
Figure 1: Architecture of VisionMask. In this example $a=1$ so $m_{1}$ is the active mask $m_{a}$.
@@ 3.1 Problem Formulation
We approach the task of VisionMask as an imitation learning problem based on Behavioral Cloning (BC) (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011). Formally, we define the environment as a Markov Decision Process (MDP) $\{\mathcal{S},\mathcal{A},P,R,\gamma\}$, where $\mathcal{S}$ represents the state space; $\mathcal{A}$ denotes the action space; the state transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ depicts the transition between states based on actions, where $\Delta(\mathcal{S})$ represents the set of probability distributions over states; the reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides the immediate reward for state-action pairs; and $\gamma\in[0,1]$ and $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ represent the discount factor and policy. Return $G$ is defined as $G=\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}$, and the expected cumulative reward of a policy $\pi$ is $\mathbb{E}_{\pi}[G]=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}]$, where the expectation is taken with respect to the initial state distribution, transition probabilities, and action probabilities determined by $\pi$.
VisionMask operates on a given pre-trained expert policy $\pi_{E}$ such that
$\pi_{E}\approx\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}[G],$
where $\pi^{*}$ is the optimal policy. We can obtain a dataset of expert demonstrations $\mathcal{D}_{E}=\{(s_{i},\pi_{E}(s_{i}))\}_{i=1}^{N}$ consisting of $N$ state-action pairs, from trajectories sampled while executing $\pi_{E}$ in the environment. Our goal is to learn an explainer $f_{\theta^{*}}$ that minimizes the loss
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{(s,a)\in\mathcal{D}_{E}}L(a,s,\theta),$
where $L$ is the training loss function to be discussed in section 3.3. The explainer function $f_{\theta}:\mathcal{S}\to[0,1]^{K\times d_{s}}$ (with $d_{s}$ representing the state dimension and $K$ the number of candidate actions) provides the attributions of each action to each feature in the state $s$. In the case of visual DRL, $s$ is the visual input of the agent and $d_{s}$ is the number of pixels in $s$, i.e., $d_{s}=W\times{H}$, with $W$ and $H$ representing the width and height of the visual input. The value of each element in the attribution is bounded within the range $[0,1]$, with the $(i,j)$th element indicates the $j$th pixel’s attribution to the $i$th action. This attribution will then be overlaid onto the visual input, masking unimportant pixels while retaining important ones. The overlaid visual input will be analyzed by the policy model again. We will update parameters $\theta$ of explainer $f$ so that the overlaid visual input always produces the same action as the original one. The expert policy $\pi_{E}$ remains frozen during the training process of $f_{\theta}$, meaning its parameters are fixed.

### v2

# Why the Agent Made that Decision: Contrastive Explanation Learning for Reinforcement Learning

@@ Abstract
Reinforcement learning (RL) has demonstrated remarkable success in solving complex decision-making problems, yet its adoption in critical domains is hindered by the lack of interpretability in its decision-making processes. Existing explainable AI (xAI) approaches often fail to provide meaningful explanations for RL agents, particularly because they overlook the contrastive nature of human reasoning—answering ”why this action instead of that one?” To address this gap, we propose a novel framework of contrastive learning to explain RL selected actions, named VisionMask. VisionMask is trained to generate explanations by explicitly contrasting the agent’s chosen action with alternative actions in a given state using a self-supervised manner. We demonstrate the efficacy of our method through experiments across diverse RL environments, evaluating it in terms of faithfulness, robustness and complexity. Our results show that VisionMask significantly improves human understanding of agent behavior while maintaining accuracy and fidelity. Furthermore, we present examples illustrating how VisionMask can be used for counterfactual analysis. This work bridges the gap between RL and xAI, paving the way for safer and more interpretable RL systems.
@@ 1 Introduction
Deep Reinforcement Learning (DRL) is a powerful technology in machine intelligence, widely used for many applications Sutton et al. (1999). However, understanding a DRL agent’s decision-making process is challenging, due to the inherent lack of explainability in the high-dimensional, non-linear structure of its underlying Deep Neural Network (DNN) Heuillet et al. (2021); Hickling et al. (2023). The lack of transparency undermines users’ trust, driving the development of Explainable AI (xAI).
Various methods in computer vision have been proposed to enhance the transparency of AI systems Ribeiro et al. (2016); Selvaraju et al. (2017); Bach et al. (2015); Shrikumar et al. (2017); Lundberg and Lee (2017). At the core, they share a common foundation: attributing the classifier’s outputs to more interpretable features and using a saliency map to visualize these attributions. Their only differences are how these attributions are calculated. A high-quality attribution-based explanation should meet several key criteria. First, it should demonstrate faithfulness, meaning that including features with high attribution should lead the model to the target output, and excluding them should prevent it. Second, it should exhibit specificity, ensuring that only critical features receive high attribution. Sometime this is also referred to as sparseness. Finally, it should be robust, meaning the explanation should remain consistent and not change significantly with minor variations in the input.
Attribution-based explanation has been studied for DRL models. Greydanus et al. (2018) and  Puri et al. (2020) utilized policy distributional shifts as the basis for attribution in RL. Specifically, they calculate attribution of a feature as the difference in Q/V values or action distributions between the original and perturbed states. For example, given agent policy $\pi$, the attribution of a feature is proportional to $E_{s^{\prime}}(|\pi(s)-\pi(s^{\prime})|_{2})$ where $s$ stands for the original state and $s^{\prime}$ represents perturbed states generated for this feature. By calculating the attributions for all features, a saliency map $m$ can be created. However, the perturbation-based explanations lack faithfulness. Since each perturbation focuses only on local features while ignoring the joint impact of feature combinations, overlaying the saliency map with the original state, $(m\odot s)$, does not result in a feature combination that leads the agent to the target action distribution $\pi(s)$.
A better approach to enhance faithfulness is to learn a model to predict the saliency map $m$ that minimizes the difference between $\pi(m\odot s)$ and $\pi(s)$. Explainer Stalder et al. (2022) leveraged this idea by training an explanation model for an image classifier. However, Explainer categorizes class labels into target and non-target for each training sample and focus on learning saliency map (or mask) only for the target label while treat all non-target labels as a single group. Unlike a (well trained) image classifier, where predictions for non-target labels are typically close to 0, DRL agent in many scenarios, does not exhibit a clear preference for the actions. Non-target actions may sometimes have probabilities only slightly lower than those of target actions. Analyzing how masking the feature may affect the non-target action probability provides additional information that can be used to train the explanation model more effectively.
The above analysis motivates us to design a trainable saliency map generator for attribution-based explanations and train it using two channels of contrastive information:(i) Action-wise contrast: We believe that environment states contain features that motivate the DRL agent to select both target action and non-target actions. However, the target action is ultimately chosen because it corresponds to higher reward or has a stronger presence. For each action, a saliency map can be generated as an explanation. Choosing features according to the saliency map for a non-target action should push the agent away from the target action, and vice versa. This inspired us to treat the saliency map of the non-target action $m_{a^{\prime}}$ and the target action $m_{a}$ as a negative pair, which can be leveraged for contrastive learning Chopra et al. (2005); Schroff et al. (2015); Gutmann and Hyvärinen (2010). (ii) Feature-wise contrast: To exclude irrelevant features (e.g. background) from the saliency map, explanations also need to be discriminative in filtering out such information. When only irrelevant features are accessible to the agent, the resulting action distribution should be as uniform as possible. Therefore, the target action’s saliency map ($m$) and its inverted counterpart ($\tilde{m}=1-m$) form another negative pair for contrastive learning.
In this work, we present VisionMask as an RL explainer that is contrastively trained to generate saliency maps to explain agent’s actions. We specifically focus on agents that maps images to actions and consider each pixel value as the interpretable input feature, although the similar technique could be extended to other type of features. We carefully design the objective function to enable self-supervised contrastive learning of explanations from both action-wise and feature-wise perspectives, fostering the generation of more faithful explanations. We conduct evaluation on six RL environments with five baselines based on faithfulness, robustness, and sparseness. Quantitatively, VisionMask outperforms the baselines in terms of faithfulness, while exhibiting strong robustness and high sparseness. Qualitatively, we compared VisionMask with the baselines in two settings: visual comparison and human studies. In the visual comparison, VisionMask provides sharper explanations that align more closely with human interpretations, as demonstrated by counterfactual examples. In the human studies, VisionMask’s explanations help users better understand the agent’s decisions and calibrate appropriate trust.
@@
  • We present VisionMask, a novel attribution method for explainable visual reinforcement learning that generates action-specific saliency maps. To the best of our knowledge, this is the first work that learns saliency map in a self-supervised contrastive learning manner, yielding highly faithful explanations across various RL environments.
  • VisionMask is model-agnostic in that it works with any vision-based DRL agent and requires only the agent’s input state and output action during inference time. VisionMask does not make any modification on the agent, however, each VisionMask model is trained specifically for a given agent by including the fixed agent in the back-propagation path for gradient descent.
  • We conducted extensive experiments, both qualitatively and quantitatively, to demonstrate the improvements of the saliency map in terms of faithfulness, robustness, and sparseness, as well as a comprehensive ablation study for each component.
@@ 2 Background and Related Work
The nested structure and non-linear operation of DNN make it challenging for humans to understand how the outputs are derived from inputs. Some existing works address this challenge by explaining model outputs through input attributions Petsiuk et al. (2018); Ribeiro et al. (2016); Fong and Vedaldi (2017). For example, Randomized Input Sampling for Explanation of Black-box Models (RISE) Petsiuk et al. (2018) is a perturbation-based approach that explains neural image classifiers by applying randomly generated masks to the input image and assessing their impact. Local Interpretable Model-agnostic Explanations (LIME)Ribeiro et al. (2016) tries to explain local instances by approximating them within a nearby vicinity using a linear decision model, where the explainable elements are superpixels (i.e., small image regions of similar pixels).
Built on top of the DNNs, the lack of transparency of the DRL agents undermines humans trust and hinders their adoption. Several studies have addressed this challenging problems. According to Qing et al. (2022), existing efforts can be categorized into four main approaches: model explaining, reward explaining, task explaining, and state explaining.
Model explaining relies on inherently interpretable model architectures or auxiliary reasoning mechanisms to generate explanations Topin et al. (2021). However, these self-explanatory or symbolic models often suffer from decreased performance compared to state-of-the-art neural network based RL policies and may lack the representational power needed to learn more complex policies. Reward explanation typically involves using explicitly designed explainable reward functions to generate explanations Ashwood et al. (2022). Task explanation considers a policy as a composition of sub-task and action sequences, explaining the behavior in terms of relationships among sub-tasks Shu et al. (2017). These approaches often assume the existence of a reward decomposition scheme or a predefined sub-task partition mechanism, which may not hold true in all RL environments.
@@
State explaining methods can further be divided into three categories: attention-based, Shapley value-based and perturbation-based mechanisms, which are detailed below. Attention is a common approach used for explainable RL in several existing works Annasamy and Sycara (2019). However, similar to i-DQN Annasamy and Sycara (2019), these methods cannot explain a given DRL model because their attention model must be trained alongside the agent model. Shapley value Shapley and others (1953), a concept from game theory, has also been introduced into XRL. SVERL Beechey et al. (2023) provides a theoretical analysis on how Shapley values can be applied to explain value and policy network in DRL. However, this approach is still in its early stages and assumes prior knowledge of the transition model of the environment, which is not realistic for realistic applications.
Perturbation-based methods Greydanus et al. (2018); Iyer et al. (2018); Puri et al. (2020) compute saliency maps of the input features by comparing the action probability or value function before and after perturbation. Greydanus et al. (2018) perturbed the state with Gaussian blur at fixed intervals, while Iyer et al. (2018) identified objects and perturbed them with a constant gray value. Puri et al. (2020) improved on Greydanus et al. (2018) by removing the effect of perturbations from other irrelevant actions. As discussed in section 1, the performance of these techniques is constrained by predefined perturbation rules and the lack of knowledge accumulation. Furthermore, there is no guarantee that the perturbed input remains physically meaningful.
Figure 1: Architecture of VisionMask. State-action pairs $(s,a)$ are sampled from $\mathcal{D}_{E}$.
@@ 3 VisionMask
In this section, we present our VisionMask architecture. The primary goal is to generate action-wise saliency maps that attribute the most relevant features in the state to each action. For agents that map images to actions, the features and states correspond to pixels and images.
@@ 3.1 Problem Formulation
Formally, we define the environment as a Markov Decision Process (MDP) $\{\mathcal{S},\mathcal{A},P,R,\gamma\}$, where $\mathcal{S}$ represents the state space; $\mathcal{A}$ denotes the action space with $|\mathcal{A}|=K$; the state transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ depicts the transition between states based on actions, where $\Delta(\mathcal{S})$ represents the set of probability distributions over states; the reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides the immediate reward for state-action pairs; $\gamma\in[0,1]$ is discount factor and $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ represent policy. Return $G$ is defined as $G=\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}$, and the expected cumulative reward of a policy $\pi$ is $\mathbb{E}_{\pi}[G]=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}]$, where the expectation is taken with respect to the initial state distribution, transition probabilities, and action probabilities determined by $\pi$. VisionMask operates on a given trained expert policy $\pi_{E}$ such that $\pi_{E}\approx\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\pi}[G],$ where $\pi^{*}$ is the optimal policy. We can obtain a dataset of expert demonstrations $\mathcal{D}_{E}=\{(s_{i},a_{i}\sim\pi_{E}(s_{i}))\}_{i=1}^{N}$ consisting of $N$ state-action pairs, from trajectories sampled while executing $\pi_{E}$ in the environment. Our goal is to learn an explainer $f_{\theta^{*}}$ that minimizes the loss $\theta^{*}=\arg\min_{\theta}\sum_{(s,a)\in\mathcal{D}_{E}}\mathcal{L}(a,s,\theta)$ where $\mathcal{L}$ is the training loss function to be discussed in section 3.3. The explainer function $f_{\theta}:\mathcal{S}\to[0,1]^{K\times d_{s}}$, where $d_{s}$ represents the feature size of the state $s\in\mathcal{S}$ and $K$ denotes the number of candidate actions, predicts the attributions of each action to each feature in the state $s$. The value of the output is bounded within the range $[0,1]$, with the $(i,j)$th element indicating the $j$th feature’s attribution to the $i$th action.
In the case of visual DRL, $s$ is the visual input of the agent and $d_{s}$ is the number of pixels in $s$, i.e., $d_{s}=W\times{H}$, with $W$ and $H$ representing the width and height of the visual input. The output of the explainer can be viewed as $K$ saliency maps, $M=\{m_{0},m_{1},....,m_{K-1}\}$, each corresponds to an action. A saliency value $m_{i}[x,y]$ near 1 indicates that the pixel $(x,y)$ has significant contributions to action $i$, whereas a value near 0 denotes irrelevance. Overlaying the $i$th saliency map with the state $s$ will highlight the input features that lead the agent to the $i$th action.